y=√(Secx- tanx/sec x+ tanx) .Prove that dy/dx= sec x ( tanx+ secx)

Grade:11

2 Answers

Vijay Mukati
7 years ago
Hint:
Here you need to apply the chain rule.
Step 1:Differentiation of y = root ( f(x)) = 1/(2*root(f(x) * diff of f(x).
Step 2:Then again for diff of f(x), you will need to apply the u/v form of differentiation.

Thanks, Vj
Nandana
7 years ago
I think it’s not sec (x) * [tan (x) + sec (x) ] !
     it is sec (x) * (tan (x) -sec (x) )
coming to this solution ,
 Given ,
  y=√(Secx- tanx/sec x+ tanx)
   divide numerator with sec x – tan x & multiply numerator with sec x – tan x , then  we get
        y = sec x – tan x
        y’ = sec x . tan x  – sec2 x = sec (x) [ tan x – sec x ] is the solution for above problem .
Thank you & check once & please inform me if  I make any mistake . . .